suppose angle a is a complement of angle b. Find the value of x, angle a and angle b is angle a= 7x + 9 and angle b= 4x + 9
Found 2 solutions by Boreal, Fombitz:
Answer by Boreal(15235)

Complements add to 90
7x+9+4x+9=90
11x=72
x=(72/11) degrees (about 6.55)
7x=504/11
9=99/11
Angle a=603/11 or 54.82 degrees
-------------
Angle b=288/11 +99/11 or 35.18 degrees

Answer by Fombitz(32388)